OpenStudy (anonymous):

I don't understand how: *****The question is in the comments made by me I used the equation writer (SORRY)**** I assume there is a rule that is missing but which one? ----I commented below so you can see what it is supposed to look like with the equation writer.-----

OpenStudy (jdoe0001):

.... wats dat?

OpenStudy (jdoe0001):

you could use the whiteboard maybe

OpenStudy (anonymous):

\[B _{2}\cos(\frac{ \pi n x }{ a })\] and \[B _{1}\sin(\frac{ \pi n x }{ a })\] equal \[Bsin(\frac{ \pi n }{ a } x + \frac{ \pi n }{ 2 })\]

OpenStudy (jdoe0001):

hmmm.... is probably easier if ... you post a quick screenshot of it

OpenStudy (jdoe0001):

maybe it's just me.... but can't see what it's meant to do

OpenStudy (anonymous):

It says it is convenient to write both cases using a single formula.